OpenStudy (anonymous):

Problem set 1c-2 Let f(x)=(x − a)g(x). Use the definition of the derivative to calculate that f'(a)= g(a), assuming that g is continuous. In the solution how does f(a) go to 0?

OpenStudy (anonymous):

in problem set 1.

OpenStudy (shubha):

f(x) = (x - a) g(x) so when x = a then f(a) = (a - a) g(a) = 0 (we substituted x = a in the function so (x - a) becomes (a - a) which is 0)

OpenStudy (anonymous):

Hi, Please find attached my proposed solution.